# Caching-Augmented Lifelong Multi-Agent Path Finding

### v2

@@ Abstract
Multi-Agent Path Finding (MAPF), which involves finding collision-free paths for multiple robots, is crucial in various applications. Lifelong MAPF, where targets are reassigned to agents as soon as they complete their initial targets, offers a more accurate approximation of real-world warehouse planning. In this paper, we present a novel mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new type of map grid called cache for temporary item storage and replacement, and designed a locking mechanism for it to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. We identified three main factors significantly impacting CAL-MAPF performance through experimentation: suitable input task distribution, high cache hit rate, and smooth traffic. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions, maps, and agent configurations.
@@ I INTRODUCTION
Automated warehouses represent a multibillion-dollar industry led by corporations such as Amazon, Alibaba, Ocado, and inVia. These facilities deploy hundreds of robots to transport goods from one location to another [1]. Planning collision-free solutions for hundreds of robots is a key component of these automated warehouses, and can be simplified to a multi-agent path finding problem (MAPF) [2]. The MAPF problem requires planning collision-free paths for multiple agents from their start locations to pre-assigned target locations in a known environment while minimizing a given cost function. Many algorithms have been developed to solve this problem optimally and suboptimally, such as Conflict Based Search (CBS) [3], $M^{*}$[4] and Enhanced CBS (ECBS) [5].
@@
Although MAPF is classified as an NP-hard problem [6], it still represents a simplistic approximation of actual warehouse planning. It is the ’one-shot’ version of the real application challenge, where an agent only needs to reach one target location and remains stationary until every participant has arrived at their respective targets. To address this limitation, an advanced version called Multi-Agent Pickup and Delivery (MAPD) or Lifelong MAPF [7] has been introduced. This version reassigns targets to agents once they complete their initial objectives, better reflecting continuous operational scenarios. Various algorithms, such as RHCR [8], MAPF-LNS [9], PIBT [10], and LaCAM [11], have been created and applied in practical environments.
@@
Although many studies focus on developing effective algorithms, there is also interest in improving the throughput of automated warehouses within the Lifelong MAPF framework by optimizing warehouse layouts. Prominent approaches include Fishbone design [12], DSAGE [13], and GGO [14]. However, these works mainly focus on static storage design and do not consider input distributions with specific patterns that may change over time. To address this problem, some studies have explored intelligent decision-making strategies [15, 16, 17]. Notable examples include the Kiva system and various policies, such as the velocity-based policy [16], and mixed-integer optimization and heuristic methods [18]. However, these methods are designed for specific warehouse systems, making it difficult to apply them to different warehouse types. They also require implementing complex rearrangement policies for the entire warehouse storage.
@@
In this paper, we present a dynamic cache delivery mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), inspired by cache design, a foundational element extensively utilized in computer architecture, databases, and various other domains. This cache mechanism is adaptable to various Lifelong MAPF algorithms, warehouse storage strategies, and incoming task distributions. We have developed a novel map grid type that allows items to be temporarily stored and replaced. Additionally, we design a lock mechanism to enhance the stability of the planning solution. We have evaluated this cache mechanism using different cache replacement policies and a range of input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: input task distribution, cache hit rate, and map design. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.
@@ II Problem Definition
Lifelong Multi-Agent Path Finding (Lifelong MAPF) problem presents unique challenges due to the dynamic and unending nature of the tasks. Following the terminology in the MAPF community, we use agents to refer to robots. Let $I=\{1,2,\cdots,N\}$ denote a set of $N$ agents. $G=(V,E)$ represents an undirected graph, where each vertex $v\in V$ represents a possible location of an agent in the workspace, and each edge $e\in E$ is a unit-length edge between two vertices that moves an agent from one vertex to the other. Self-loop edges are allowed, which represent “wait-in-place” actions. Each agent $i\in I$ has a unique start location $s_{i}\in V$. Assume there is an task queue $Q=[q_{1},q_{2},...]$. For each task $q_{i}\in Q$, $q_{i}$ represents a certain item located in a reachable vertex in $V$ where at least one agent can find a path to the vertex. Each item must be delivered to a vertex in $V$ as the user specifies. These tasks are generated from a specific probability distribution to enhance realism.
@@
Figure 1: Caching-Augmented Map: (1) Blue grids represent Shelves $S$. (2) Purple grids represent Caches $C$. (3) Green grids represent unloading ports $U$. The upper map has multiple ports, while a single port is in the bottom one. In the multi-port map, each unloading port has an independent cache area, task queue, and agents. The cache areas are near the unloading ports within ±$\pm 2$ rows. For the single-port map, the port can utilize all agents and all caches. Given that the number of cache grids can affect the cache hit rate, we also tested different numbers of cache grids, ranging from 80 to 16, by removing cache grids column by column from right to left.
@@
We study an on-line setting in which incoming tasks are not known in advance. It is assumed that there is a task assigner (external to low-level MAPF algorithm) from which agents can request target locations throughout the system’s operation. The task assigner determines the specific target locations for agents based on tasks in $Q$. In our paper, this task assigner is considered naive and will return the target location based on the first unfinished task in $Q$. Each agent $i\in I$ starts from a start location $s_{i}\in V$ and its target location depends on the task assigner. As agents complete their paths at their assigned goals, they are immediately assigned new targets, reflecting warehouse operations’ perpetual and iterative nature. Each action of agents, waiting in place or moving to an adjacent vertex, takes a unit of time. An agent’s path, $p^{i}=[v_{0}^{i},v_{1}^{i},\ldots,v_{T_{i}}^{i}]$, tracks the sequence of vertices traversed by agent $i$ from its start to its target.
@@
As shown in  Figure 1, we primarily focus on 2D warehouse layout maps and categorize all non-obstacle map grids into three types: blue grids $S=\{s_{i}|s_{i}\in V\}$ for shelves that store items, white grids $W=\{w_{i}|w_{i}\in V\}$ for normal aisle grids, and green grids $U=\{u_{i}|u_{i}\in V\}$ for unloading ports where agents deliver items to complete a task. Agents can enter $S$ positions only from $W$ positions. Direct transitions between two positions in $S$ are not allowed.
Next, we will introduce specific assumptions in this paper beyond the standard Lifelong MAPF definition. We assume each shelf $S$ stores an infinite number of only one unique type of item. There will be a total of $M$ different types of items corresponding to the number of shelves. Each unloading port $u_{k}$ is associated with an independent task queue $Q_{k}$. Each task $q_{i}\in Q_{k}$ must be delivered to the unloading port $U_{k}$.
We will also involve a new type of map grid: caches. These additional vertices, designated as $C=\{c_{i}|c_{i}\in V\}$ in purple, serve as interim storage areas to streamline the completion of tasks by reducing the travel time of the agents retrieving items. To the best of our knowledge, no previous MAPF works have explored this new cache mechanism, so we adopt the following simplifying assumptions: (1) The warehouse is considered to have an unlimited supply of items, eliminating worries about stock running out. (2) Agents have infinite capacity to carry items. However, they are limited to transporting one type of item at a time, enabling them to move any quantity of that specific item without restriction. (3) Items removed or evicted from the cache are treated as if they vanish immediately.
@@ III-A MAPF
(One-Shot) MAPF, which has been proved an NP-hard problem with optimality [6], has a long history [19]. It has inspired a wide range of solutions for its related challenges. Decoupled strategies, as outlined in [19, 20, 21], approach the problem by independently planning paths for each agent before integrating these paths. In contrast, coupled approaches [22, 23] devise a unified plan for all agents simultaneously. There also exist dynamically coupled methods [3, 24] that consider agents that are planned independently at first and then together only when needed in order to resolve agent-agent conflicts. Among these, Conflict-Based Search (CBS) algorithm [3] stands out as a centralized, optimal solution for MAPF, with several bounded-suboptimal variants such as ECBS [5] and EECBS [25]. Some suboptimal MAPF solvers, such as Prioritized Planning (PP) [26, 19], PBS [27], PIBT [10] and their variant methods [28, 29, 11, 30] exhibit better scalability and efficiency.
@@ III-B Lifelong MAPF
Compared to the MAPF problem, Lifelong MAPF continuously assigns new target locations to agents once they have reached their current targets. It is not necessary for agents to arrive at their targets simultaneously in Lifelong MAPF. There are mainly three kinds of methods: solving the problem as a whole [31], using MAPF methods but replanning all paths at each specified length timestep [8, 10], and replanning only when agents reach their current target locations and are assigned new targets [32, 33, 11, 30]. Among them, MAPF-LNS [9, 29] and LaCAM [11, 30] are the leading methods with high performance in task execution.
@@ III-C Cache
The cache [34] serves as a crucial hardware component designed to temporarily store data, enhancing the efficiency of future request processing. Its main goal is to speed up access to frequently used data, thus minimizing dependence on slower storage disks. Popular caching policies include Least Recently Used (LRU) [35] and First-In-First-Out (FIFO) [36], with studies indicating LRU’s superiority over FIFO [37, 38, 39]. The caching concept is widely applied in various fields, such as Databases [40] and Content Delivery Networks (CDN) [41].
@@ III-D Warehouse Storage Strategy
Automated Storage and Retrieval Systems (AS/RS) have attracted attention for their potential to enhance warehouse efficiency and reduce operational costs [42, 43]. Various strategies for assigning items to storage locations have been widely adopted and evaluated [43, 44]. The random storage policy allocates each item type to a randomly chosen storage location and offers high space utilization [45]. The policy for storage at the closest open location places the arriving unit of items at the nearest available storage location, minimizing the immediate travel distance [46, 47]. The turnover-based storage policy assigns storage locations to each item based on their demand frequency [16, 17].
@@ IV Method
In this section, we introduce our novel Caching-Augmented Lifelong MAPF (CAL-MAPF) framework along with a cache lock mechanism. Consider a warehouse scenario as depicted in Figure 1. When the task assigner encounters a new task requiring completion, the first step typically involves determining which location should be allocated to an available agent. The task assigner first checks whether the task item is already stored in the cache grids. If located, the task assigner will direct agents to retrieve the item from the cache grid and deliver it to the unloading port, or alternatively, the agent may need to collect the item from the shelves. Subsequently, the item is stored in a cache grid. Once stored, the agent can transport the item from the cache to an unloading port.
However, similar to caching in computer architecture, there is a risk that other agents could replace an item in the cache before it being retrieved by assigned agents in multi-agent scenarios. This situation could result in a low-level MAPF algorithm, such as LaCAM, generating unusual collision-free plans at each timestep, causing agents to move back and forth between two adjacent grids. Thus, a lock mechanism ensures agents can secure an item after confirming its availability in cache grids.
To ensure efficient and synchronized access to cache locations in the CAL-MAPF framework, we implement a cache lock mechanism supported by a state machine. This approach effectively prevents race conditions during cache interactions. The mechanism uses read locks for agents fetching items and writes locks for agents updating the cache, with the state machine managing transitions to maintain access order and integrity. Every cache grid will have its independent read lock and write lock. This streamlined system guarantees that agents can safely access and modify cache contents, facilitating seamless task execution within the dynamic warehouse environment.
IV-A Cache Lock Mechanism
IV-A1 Lock Types
@@
IV-A2 Lock Acquisition and Release
The mechanism defines a protocol for lock acquisition and release: (1) Acquisition: To obtain a read lock, an agent must verify that no write lock is active on the desired cache location. Conversely, to secure a write lock, an agent must confirm that the target cache location is free from any active read or write locks. (2) Release: Agents release read locks after successfully retrieving the goods. Agents release write locks once they have updated the cache, making the new items available to others.
@@ IV-A2 Lock Acquisition and Release
34:    if targetLoc == portLoc then
35:         agent.status = statusa
36:    else
37:         agent.status = statusb
38:    agent.targetLoc=targetLoc
@@
IV-B Algorithm
As shown in Algorithms 1, 2 and 3, the whole procedure is mainly based on the task assigner, who encounters new tasks and decides what locations the agents need to visit. As illustrated in Figure 1, CAL-MAPF could have multiple unloading ports and we call each unloading port and cache around the port a group. Each group is independent in task queue, cache grids, and agents. Agents in one group can only accept tasks from this group and use cache grids of this group. Algorithm 1 illustrates the operation of a group. Task assigner uses a state machine to control agent status and how they transfer to another based on different situations. When any agent reaches its target location, we must check all agents, update their status, and assign new target locations.
The state machine is illustrated in Figure 2, and the pseudocode of Algorithm 1 outlines how an agent transitions through five states while performing the low-level MAPF algorithm for a group. The task assigner monitors and updates the status of agents based on the item’s location, the state of the cache, and the agents’ current status. This algorithm ensures the efficient completion of all tasks by orchestrating the agents’ paths and their interactions with the cache and unloading port.
Figure 3: The average frequency of agent wait actions on each map grid with 256 agents under Zhang distribution. It is evident that both CAL-MAPF and baseline experience severe traffic congestion as agent number is large enough.
In Algorithm 1, we first set the status of each agent to 0 and assign them initial target locations (Lines 1-6). Next, we obtain collision-free paths for each agent by invoking a low-level MAPF algorithm. The plan will be executed until at least one agent reaches its target location (Lines 7-9). Upon arrival, the task assigner updates the lock information, attempts to assign new target locations to available agents, and updates the status of all agents based on the current warehouse situation (Lines 10-11).
Task assigner should first attempt to release locks to allow other agents the opportunity to access cache grids. Status 1 is trying to read a cache grid, and Status 2 is going to write. Only agents in status 1 and status 2 can hold locks. After reaching their target location, they must go to the unloading port to deliver items. So, we verify whether these agents have reached their targets before releasing their held locks (Algorithm 2 Lines 1-8).

### v3

@@ Abstract
Multi-Agent Path Finding (MAPF), which involves finding collision-free paths for multiple robots, is crucial in various applications. Lifelong MAPF, where targets are reassigned to agents as soon as they complete their initial targets, offers a more accurate approximation of real-world warehouse planning. In this paper, we present a novel mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new type of map grid called cache for temporary item storage and replacement, and created a locking mechanism to improve the planning solution’s stability. A task assigner (TA) is designed for CAL-MAPF to allocate target locations to agents and control agent status in different situations. CAL-MAPF has been evaluated using various cache replacement policies and input task distributions. We have identified three main factors significantly impacting CAL-MAPF performance through experimentation: suitable input task distribution, high cache hit rate, and smooth traffic. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions, map and agent configurations.
@@ I INTRODUCTION
Automated warehouses represent a multibillion-dollar industry led by corporations such as Amazon, Ocado, and inVia. These facilities deploy hundreds of robots to transport goods from one location to another [1]. Planning collision-free solutions for hundreds of robots is a key component of these automated warehouses, and can be simplified to a multi-agent path finding problem (MAPF) [2]. The MAPF problem requires planning collision-free paths for multiple agents from their start locations to pre-assigned target locations in a known environment while minimizing a given cost function. Many algorithms have been developed to solve this problem optimally and suboptimally, such as Conflict Based Search (CBS) [3], $M^{*}$[4] and Enhanced CBS (ECBS) [5].
@@
Although MAPF is classified as an NP-hard problem [6], it still represents a simplistic approximation of actual warehouse planning. It is the ’one-shot’ version of the real application challenge, where an agent only needs to reach one target location and remains stationary until every agent has arrived at their respective targets. To address this limitation, a new problem called Multi-Agent Pickup and Delivery (MAPD) or Lifelong MAPF [7] has been introduced. This version assigns new targets to agents once they reach their current targets, reflecting continuous operational scenarios better. Various algorithms, such as RHCR [8], MAPF-LNS [9], PIBT [10], and LaCAM [11], have been created and applied in practical environments.
@@
Although many studies focus on developing effective algorithms, there is also interest in improving the throughput of automated warehouses within the Lifelong MAPF framework by optimizing warehouse layouts. Prominent approaches include Fishbone design [12], DSAGE [13], and GGO [14]. However, these works mainly focus on static storage design and do not consider input distributions with specific patterns that may change over time. To address this problem, some studies have explored intelligent decision-making strategies [15, 16, 17]. Notable examples include Kiva system and various policies, such as the velocity-based policy [16], and mixed-integer optimization and heuristic methods [18]. However, these methods are designed for specific warehouse systems, making applying them to different warehouses difficult. They also require implementing complex rearrangement policies for the entire storage.
@@
In this paper, we present a dynamic cache delivery mechanism named Caching-Augmented Lifelong MAPF (CAL-MAPF), inspired by cache design, a foundational idea extensively utilized in computer architecture, databases, and various other domains. This cache mechanism is adaptable to multiple Lifelong MAPF algorithms, warehouse storage strategies, and incoming task distributions. We developed a new map grid type that allows items to be temporarily stored and replaced. We also designed a lock mechanism to enhance the stability of the planning solution and a task assigner (TA), which allocates target locations to agents and controls agent status in different situations. We have evaluated this cache mechanism using different cache replacement policies and input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: input task distribution, cache hit rate, and map design. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions, map and agent configurations.
@@ II Problem Definition
Lifelong Multi-Agent Path Finding (Lifelong MAPF) problem presents unique challenges due to the dynamic and unending nature of the tasks. Let $I=\{1,2,\cdots,N\}$ denote a set of $N$ agents. $G=(V,E)$ represents an undirected graph, where each vertex $v\in V$ represents a possible location of an agent in the workspace, and each edge $e\in E$ is a unit-length edge between two vertices that moves an agent from one vertex to the other. Self-loop edges are allowed, which represent “wait-in-place” actions. Each agent $i\in I$ has a unique start location $s_{i}\in V$. Assume there is an task queue $Q=[q_{1},q_{2},...]$. For each task $q_{i}\in Q$, $q_{i}=(idx,loc)$ represents a certain kind of item with type number $idx$ should be sent to position $loc$. This $idx$ item should be located in a reachable vertex in $V$ where at least one agent can find a path to the vertex. Each item must be delivered to a vertex in $V$ as the user specifies. These tasks are generated from a specific probability distribution to enhance realism.
@@
Figure 1: Caching-Augmented Maps: (1) Blue grids represent Shelves $S$. (2) Purple grids represent Caches $C$. (3) Green grids represent unloading ports $U$. The upper map has multiple ports, while a single port is in the bottom one. In the multi-port map, each unloading port has an independent cache area, task queue, and agents. The cache areas are near the unloading ports within ±$\pm 2$ rows. For the single-port map, the port can utilize all agents and all caches. Given that the number of cache grids can affect the cache hit rate, we also tested different numbers of cache grids, ranging from 80 to 16, by removing cache grids column by column from right to left.
@@
We study an online setting in which incoming tasks are not known in advance. It is assumed that there is a task assigner (external to MAPF algorithm). The TA determines the specific target locations for agents based on tasks in $Q$. In our paper, TA is considered naive and will return the target location based on the first unassigned task in $Q$ for each spare agent. Each agent $i\in I$ starts from a location $s_{i}\in V$ and its target location depends on the TA. Each action of agents, which could be waiting in place or moving to an adjacent vertex, takes a unit of time. An agent’s path, $p^{i}=[v_{0}^{i},v_{1}^{i},\ldots,v_{T_{i}}^{i}]$, tracks the sequence of vertices traversed by agent $i$ from its start to its target. As agents complete their paths at their assigned targets, they are immediately assigned new target locations, reflecting warehouse operations’ perpetual and iterative nature.
@@
As shown in  Figure 1, we primarily focus on 2D warehouse layout maps and categorize all non-obstacle map grids into three types: blue grids $B=\{b_{i}|b_{i}\in V\}$ for shelves that store items, white grids $W=\{w_{i}|w_{i}\in V\}$ for normal aisle grids, and green grids $U=\{u_{i}|u_{i}\in V\}$ for unloading ports where agents deliver items to complete a task. Agents can enter $B$ positions only from $W$ positions. Direct transitions between two positions in $B$ are not allowed.
@@ III-A MAPF
(One-Shot) MAPF, which has been proved an NP-hard problem with optimality [6], has a long history [19]. This problem is finding collision-free paths for multiple agents while minimizing a given cost function. It has inspired a wide range of solutions for its related challenges. Decoupled strategies, as outlined in [19, 20, 21], approach the problem by independently planning paths for each agent before integrating these paths. In contrast, coupled approaches [22, 23] devise a unified plan for all agents simultaneously. There also exist dynamically coupled methods [3, 24] that consider agents planned independently at first and then together only when needed in order to resolve agent-agent conflicts. Among these, Conflict-Based Search (CBS) algorithm [3] stands out as a centralized and optimal method for MAPF, with several bounded-suboptimal variants such as ECBS [5] and EECBS [25]. Some suboptimal MAPF solvers, such as Prioritized Planning (PP) [26, 19], PBS [27], PIBT [10] and their variant methods [28, 29, 11, 30] exhibit better scalability and efficiency.
@@ III-B Lifelong MAPF
Compared to the MAPF problem, Lifelong MAPF continuously assigns new target locations to agents once they have reached their current targets. Agents don’t need to arrive at their targets simultaneously in Lifelong MAPF. There are mainly three kinds of methods: solving the problem as a whole [31], using MAPF methods but replanning all paths at each specified length timestep [8, 10], and replanning only when agents reach their current target locations and are assigned new targets [32, 33, 11, 30]. Among them, MAPF-LNS [9, 29] and LaCAM [11, 30] are the leading methods.
@@ III-C Cache
The cache [34] serves as a crucial component in computer science designed to temporarily store data, enhancing the efficiency of future request processing. Its main goal is to speed up access to frequently used data, thus minimizing dependence on slower storage disks. Popular caching policies include Least Recently Used (LRU) [35] and First-In-First-Out (FIFO) [36], with studies indicating LRU’s superiority over FIFO [37, 38, 39]. The caching concept is widely applied in various fields, such as Databases [40] and CDN [41].
@@ III-D Warehouse Storage Strategy
Automated Storage and Retrieval Systems (AS/RS) have attracted attention for their potential to enhance warehouse efficiency and reduce operational costs [42, 43]. Various strategies for assigning items to storage locations have been widely adopted and evaluated [43, 44]. The random storage policy allocates each item type to a randomly chosen storage location and offers high space utilization [45]. The closest storage policy places new items at the nearest available storage location to minimize the immediate travel distance [46, 47]. The turnover-based storage policy assigns items to storage locations based on their demand frequency [16, 17].
@@ IV Method
In this section, we introduce our Caching-Augmented Lifelong MAPF (CAL-MAPF) framework along with a new map grid design, a TA and a cache lock mechanism.
IV-A Cache Grid
As shown in Figure 1, we assume each shelf $b_{i}\in B$ stores an infinite number of only one unique type of item. There will be a total of $M$ different types of items corresponding to the number of shelves. Each unloading port $u_{k}$ is associated with an independent task queue $Q_{k}$. Each task $q_{i}\in Q_{k}$ must be delivered to the unloading port $U_{k}$. We will also involve a new type of map grid: caches. These additional vertices, designated as $C=\{c_{i}|c_{i}\in V\}$ in purple, serve as interim storage areas to streamline the completion of tasks by reducing the travel time of the agents retrieving items.
To the best of our knowledge, no previous MAPF works have explored this new cache mechanism, so we adopt the following simplifying assumptions: (1) The warehouse has an unlimited supply of items, eliminating worries about stock running out. (2) Agents have infinite capacity to carry items. However, they are limited to transporting one type of item at a time, enabling them to move any quantity of that specific item without restriction. (3) Items removed or evicted from the cache are treated as if they vanish immediately.
IV-B Task Assigner
The task assigner (TA) is external to MAPF algorithm but can define all agent’s target locations at any timestep. When the TA encounters a new task in $Q$ requiring completion, the first step typically determines where to get the item and which location should be allocated to an available agent. So, the TA first checks whether the task item exists in the cache grids. If so, TA will assign an agent with the cache location of the item. After the agent gets the item, TA then assigns the unloading port location to it. If not in cache grids, TA will give the item’s shelf location to the agent and then one insertable cache grid location. Subsequently, the item is stored in the cache grid. Once stored, the agent can transport the item from the cache to an unloading port.
However, similar to caching in computer architecture, there is a risk that other agents could replace the item in the cache with other items in multi-agent scenarios. This situation could result in MAPF algorithm, such as LaCAM, generating unusual collision-free plans at each timestep, causing agents to move back and forth between two adjacent grids. Thus, a lock mechanism ensures agents can secure an item after confirming its availability in cache grids.
IV-C Cache Lock Mechanism
In the CAL-MAPF framework, TA has a cache lock mechanism supported by a state machine to ensure efficient and synchronized access to cache locations. This approach effectively prevents race conditions during cache interactions. The mechanism uses read locks for agents fetching items and writes locks for agents updating the cache. Every cache grid will have its independent read lock and write lock. This streamlined system guarantees that agents can access and modify cache contents safely, facilitating seamless task execution within the dynamic warehouse environment.
IV-C1 Lock Types
@@
IV-C2 Lock Acquisition and Release
The mechanism defines a protocol for lock acquisition and release: (1) Acquisition: To obtain a read lock, an agent must verify no write lock is active on the desired cache location. Conversely, to secure a write lock, an agent must confirm that the target cache location is free from any active read or write locks. (2) Release: Agents release read locks after successfully retrieving items. Agents release write locks once they have updated the cache, making the inserted new items available to others.
@@ IV-C2 Lock Acquisition and Release
34:    agent.status = statusa
35:    if targetLoc != portLoc then
36:         agent.status = statusb
37:    agent.targetLoc=targetLoc
@@
IV-D Algorithm
As shown in Algorithms 1, 2 and 3, we will introduce the whole procedure of CAL-MAPF. It is mainly based on the TA, who encounters new tasks and decides what locations the agents need to visit. As illustrated in Figure 1, CAL-MAPF could have multiple unloading ports and we call each unloading port and cache around the port a group. Each group is independent in task queue, cache grids, and agents. Agents in one group can only accept tasks from this group and use cache grids of this group. Algorithm 1 illustrates the operation of one group.
Figure 3: The average frequency of agent wait actions on each map grid with 256 agents under Zhang distribution. As the agent number is large enough, both CAL-MAPF and baseline experience severe traffic congestion. The congestion position bias in the map could be caused by the randomization of item indexes and tasks in $Q$.
TA uses a state machine to control agent status and how they transfer to another based on different situations. When any agent reaches its target location, TA must check all agents, update their status, and assign new target locations. The state machine is illustrated in Figure 2, and the pseudocode of Algorithm 1 outlines how an agent transitions through five states while performing the MAPF algorithm for a group. The TA monitors and updates the status of agents based on the item’s location, the state of the cache, and the agents’ current status. This algorithm ensures the efficient completion of all tasks by orchestrating the agents’ paths and their interactions with the cache and unloading port.
Overall, CAL-MAPF needs to use TA to assign target locations to all spare agents based on $Q$. Then invoke MAPF solver to find a collision-free solution and execute it until at least one agent reaches its target location. Then TA will check all agent’s statuses, release locks and change agent status following the state machine. When changing status, TA needs to use tasks in $Q$ and cache replacement policy to determine target locations for agents and acquire locks for agents. CAL-MAPF repeats this process during the lifelong scenario.
In Algorithm 1, we first set the status of each agent to 0 and assign them initial target locations (Lines 1-6). Next, we obtain collision-free paths for each agent by invoking a MAPF algorithm. The plan will be executed until at least one agent reaches its target location (Lines 7-9). Upon arrival, TA updates the lock information, attempts to assign new target locations to available agents, and updates the status of all agents based on the current warehouse situation (Lines 10-11).
TA should first attempt to release locks to allow other agents the opportunity to access cache grids. Status 1 is trying to read a cache grid, and Status 2 is going to write. Only agents in status 1 and status 2 can hold locks. After reaching their target location, they must go to the unloading port to deliver items. So, we verify whether these agents have reached their targets before releasing their held locks (Algorithm 2 Lines 1-8).